# Deformations of the Fano scheme of a cubic

### v2

@@ Introduction
$\mathrm{H}^{0}(\Theta_{\mathbf{P}^{{}^{\prime}}})\xrightarrow{\sim}\mathrm{H}^{0}(\Theta_{\mathbf{P}^{{}^{\prime}}}|_{\mathrm{Z}}),\quad\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}_{d}}(2))/(h)\xrightarrow{\sim}\mathrm{H}^{0}(\mathcal{O}_{\mathrm{Z}}(2)),$
@@
Since $\mathrm{X}$ is irreducible, the second equation implies $l_{d+1}=0$ and $\lambda=3\mu$, while the first one gives $2\mu+\lambda_{d+1}=\lambda$. ∎
@@
for any nodal cubic. Low-dimensional ($2\leqslant d\leqslant 4$) nodal cubics are in fact known to be stable in the sense of geometric invariant theory [18], and so this holds for $2\leqslant d\leqslant 4$.
@@
Instead of $\mathcal{D}_{\mathrm{X}}$, one could consider the subfunctor $\mathcal{D}_{\mathrm{X}}^{{}^{\prime}}$ of $\mathcal{D}_{\mathrm{X}}$ given by the locally trivial deformations of $\mathrm{X}$; here $\mathrm{H}^{2}(\Theta_{\mathrm{X}})$ is an obstruction space of $\mathcal{D}_{\mathrm{X}}^{{}^{\prime}}$ [24]. While it is known that if $d=2$ or $d=3$, then $\mathrm{H}^{2}(\Theta_{\mathrm{X}})=0$ [21], this vanishing need not hold when $d$ is large. In fact, the following holds:
@@
Indeed, $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$ and Lemma 4.1 imply that
@@
coming from the local-to-global spectral sequence; here $\mathcal{E}xt^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})$ is the structure sheaf of the singular locus. As a special case of a result of Varchenko [27],
@@
The space $\mathrm{H}^{2}(\Theta_{\mathrm{X}})$ is canonically isomorphic to $\mathrm{H}^{1}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$, where $\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}}$ is the equisingular normal sheaf of $\mathrm{X}\subset\mathbf{P}$. We can view $\mathrm{X}$ as a point of the Hilbert scheme $\mathrm{V}^{\delta}_{d}$ of cubic hypersurfaces in $\mathbf{P}$ with $\delta$ nodes (Severi scheme); $\mathrm{H}^{0}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$ and $\mathrm{H}^{1}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$ are then the tangent and obstruction spaces of $\mathrm{V}^{\delta}_{d}$ at $[\mathrm{X}]$ [13]. Proposition 4.2 naturally leads to an extension of Theorem 111 of [7].
@@
Consider a cubic $\mathrm{X}$ with finitely many singularities, and an infinitesimal deformation $\mathfrak{X}$ of $\mathrm{X}$ over an Artin ring $\Lambda$. Then $\mathfrak{X}$ is induced by a deformation $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ of $\mathrm{X}$ in $\mathbf{P}$, and $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ is a cubic (Remark 4.1). Using the induced polarisation $\mathcal{O}_{\mathfrak{X}}(1)$ of $\mathfrak{X}$ over $\Lambda$, we can consider the relative Hilbert scheme of lines $\mathrm{F}(\mathfrak{X}/\Lambda)$, which is naturally a closed subscheme of $\mathbf{G}_{\Lambda}$. The morphism
@@
is flat, because the fibre over the closed point $\mathrm{F}=\mathrm{F}(\mathrm{X})$ is a local complete intersection (Theorem 2.2). In particular, $\mathrm{F}(\mathfrak{X}/\Lambda)$ can be thought of as an infinitesimal deformation of $\mathrm{F}$ in $\mathbf{G}$ over $\Lambda$. For any morphism of local Artin rings $\Lambda\rightarrow\Lambda^{{}^{\prime}}$, we have
@@
[…]
The proof of our main theorem requires an analogue of Lemma 4.1 for $\mathrm{F}\subset\mathbf{G}$.
@@
(i) By Corollary 2.1, the spectral sequence
@@
Here $\mathrm{H}^{0}(\mathrm{S}^{3}\mathcal{S}\otimes\mathrm{S}^{3}\mathcal{S}^{\vee})$ is one-dimensional (Proposition 2.1), and it remains to combine this with the exact sequence in cohomology associated to
$0\rightarrow\mathcal{J}_{\mathrm{F}/\mathbf{G}}\otimes\mathrm{S}^{3}\mathcal{S}^{\vee}\rightarrow\mathrm{S}^{3}\mathcal{S}^{\vee}\rightarrow\mathrm{S}^{3}\mathcal{S}|_{\mathrm{F}}^{\vee}\rightarrow 0$
(ii) Similarly, by Corollary 2.1 the spectral sequence
@@
We now apply Lemma 2.4 to $\mathcal{F}=\mathrm{S}^{3}\mathcal{S}^{\vee}$ on $\mathrm{S}=\mathbf{G}$. Let $\pi:\mathcal{Z}\rightarrow\mathrm{H}^{0}(\mathrm{S}^{3}\mathcal{S}^{\vee})$ be as in Lemma 2.4, and put $\phi=\sigma^{-1}\circ\pi$, where $\sigma$ is the isomorphism of Remark 2.1.
@@
In other words, the deformation of $\mathrm{F}$ induced by $\phi$ is complete at $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. For $d\geqslant 6$, this is a theorem of Borcea [5].
@@
There is a canonical isomorphism
@@
$\operatorname{Aut}(\mathrm{X})\rightarrow\operatorname{Aut}(\mathrm{F}).$ (15)
Since $\mathrm{X}$ can be covered by lines, this morphism is injective; by [8] the image of (15) is the subgroup $\operatorname{Aut}(\mathrm{F},\mathcal{O}_{\mathrm{F}}(1))$ of automorphisms preserving the Plücker polarization. As $\mathrm{H}^{1}(\mathcal{O}_{\mathrm{F}})=0$, $\mathrm{H}^{0}(\Theta_{\mathrm{F}})$ the tangent space of $\operatorname{Aut}(\mathrm{F},\mathcal{O}_{\mathrm{F}}(1))$ at the identity. ∎
@@
${\mathrm{H}^{0}(\mathcal{N}_{\mathrm{X}/\mathbf{G}})}$${\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})}$${\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})}$${\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})}$$\scriptstyle{d\eta_{\mathcal{H}}}$$\scriptstyle{d\eta}$
of differentials induced by (14). By Lemma 2.3 (i) and Lemma 5.1 (ii), the differential
@@
by the description of section 2.2. Since $\sigma$ is an isomorphism and the restriction map
@@
is surjective by Lemma 5.1 (i), it follows that $d\eta_{\mathcal{H}}$ is surjective. It now suffices to show that
@@
associated to the conormal sequences of $\mathrm{X}\subset\mathbf{P}$ and $\mathrm{F}\subset\mathbf{G}$, respectively. By Lemma 5.2 we have $h^{0}(\Theta_{\mathrm{X}})=h^{0}(\Theta_{\mathrm{F}})$, while
@@
result from Lemma 4.1, Lemma 5.1, and Example 2.1.
If $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, then $\mathrm{H}^{0}(\Theta_{\mathrm{F}})=0$ by Lemma 5.2. Hence both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ are pro-representable; since $\mathcal{D}_{\mathrm{X}}$ is smooth and $d\eta$ bijective, it remains to apply Lemma 2.2 (ii). ∎
@@
This is a consequence of the proof of Theorem 5.1 rather than Theorem 5.1 itself. The proof shows that $d\eta_{\mathcal{H}}$ can be identified with the isomorphism
@@
induced by $\sigma$. As $\mathcal{H}_{\mathrm{X}/\mathbf{P}}$ and $\mathcal{H}_{\mathrm{F}/\mathbf{G}}$ are pro-representable, and $\mathcal{H}_{\mathrm{X}/\mathbf{P}}$ smooth, $\eta_{\mathcal{H}}$ is an isomorphism by Lemma 2.2 (ii). Finally, $\eta$ is surjective by Lemma 2.2 (i), as both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ have a pro-representable hull by Schlessinger’s theorem, Theorem 2.3 (i). ∎
@@
The proof of Theorem 5.1 depends on [8]. One could get rid of this dependence by establishing a commutative diagram
@@
where the isomorphism on the left is induced by Chow’s isomorphism $\operatorname{Aut}(\mathbf{P})\rightarrow\operatorname{Aut}(\mathbf{G})$, and Lemma 4.1 (ii), Lemma 5.1 (ii). We expect $\eta$ to be an isomorphism without assuming the condition $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$.
@@
There are number of follow-up questions. If $\mathrm{X}$ is a Lefschetz cubic with a node at $x_{0}$, then the singular locus of $\mathrm{F}$ can be identified with a smooth complete intersection $\Sigma\subset\mathbf{P}_{d}$ of type $(2,3)$. The scheme $\mathrm{F}$ has rational singularities, and the blow up
@@
of $\mathrm{F}$ along $\Sigma$ provides a resolution of singularities of $\mathrm{F}$ [9]. In such a situation, a general construction of Wahl [28] yields a blow-down morphism
@@
Here $\tilde{\mathrm{F}}$ is closely related to the Hilbert scheme of points $\Sigma^{[2]}$. By [3], one has a canonical isomorphism $\mathrm{H}^{1}(\Theta_{\Sigma})\xrightarrow{\sim}\mathrm{H}^{1}(\Theta_{\Sigma^{[2]}})$, which shows in particular that $\mathrm{H}^{1}(\Theta_{\Sigma^{[2]}})$ has dimension $\binom{d+2}{3}$; since this is also the dimension of $\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$, the morphism $\beta$ might be an isomorphism.

### v3

@@ Introduction
$\mathrm{H}^{0}(\Theta_{\mathbf{P}^{{}^{\prime}}})\xrightarrow{\sim}\mathrm{H}^{0}(\Theta_{\mathbf{P}^{{}^{\prime}}}|_{\mathrm{Z}}),\quad\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}^{{}^{\prime}}}(2))/(h)\xrightarrow{\sim}\mathrm{H}^{0}(\mathcal{O}_{\mathrm{Z}}(2)),$
@@
Since $\mathrm{X}$ is irreducible, the second equation implies $l_{d+1}=0$ and $\lambda=3\mu$, while the first one gives $\lambda_{d+1}=\lambda-2\mu=\mu$. ∎
@@
for any nodal cubic. Low-dimensional ($2\leqslant d\leqslant 4$) nodal cubics are in fact known to be stable in the sense of geometric invariant theory [17, Theorem 1.1], and so the vanishing of $\mathrm{H}^{0}(\Theta_{\mathrm{X}})$ holds for $2\leqslant d\leqslant 4$.
@@
Instead of $\mathcal{D}_{\mathrm{X}}$, one could consider the subfunctor $\mathcal{D}_{\mathrm{X}}^{{}^{\prime}}$ of $\mathcal{D}_{\mathrm{X}}$ given by the locally trivial deformations of $\mathrm{X}$; here $\mathrm{H}^{2}(\Theta_{\mathrm{X}})$ is an obstruction space of $\mathcal{D}_{\mathrm{X}}^{{}^{\prime}}$ [23, Theorem 2.4.1]. While it is known that if $d=2$ or $d=3$, then $\mathrm{H}^{2}(\Theta_{\mathrm{X}})=0$ [20, Proposition 4], this vanishing need not hold when $d$ is large. In fact, the following holds:
@@
Indeed, $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$ and Lemma˜4.1 imply that
@@
coming from the local-to-global spectral sequence; here $\mathcal{E}xt^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})$ is the structure sheaf of the singular locus. As a special case of a result of Varchenko [26, §2],
@@
The space $\mathrm{H}^{2}(\Theta_{\mathrm{X}})$ is canonically isomorphic to $\mathrm{H}^{1}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$, where $\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}}$ is the equisingular normal sheaf of $\mathrm{X}\subset\mathbf{P}$. We can view $\mathrm{X}$ as a point of the Hilbert scheme $\mathrm{V}^{\delta}_{d}$ of cubic hypersurfaces in $\mathbf{P}$ with $\delta$ nodes (Severi scheme); $\mathrm{H}^{0}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$ and $\mathrm{H}^{1}(\mathcal{N}^{{}^{\prime}}_{\mathrm{X}/\mathbf{P}})$ are then the tangent and obstruction spaces of $\mathrm{V}^{\delta}_{d}$ at $\mathrm{X}] [$12, §3]. Proposition˜4.2 naturally leads to an extension of Theorem 111 of [6].
@@
Consider a cubic $\mathrm{X}$ with finitely many singularities, and an infinitesimal deformation $\mathfrak{X}$ of $\mathrm{X}$ over an Artin ring $\Lambda$. Then $\mathfrak{X}$ is induced by a deformation $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ of $\mathrm{X}$ in $\mathbf{P}$, and $\mathfrak{X}\subset\mathbf{P}_{\Lambda}$ is a cubic (Remark˜4.1). Using the induced polarisation $\mathcal{O}_{\mathfrak{X}}(1)$ of $\mathfrak{X}$ over $\Lambda$, we can consider the relative Hilbert scheme of lines $\mathrm{F}(\mathfrak{X}/\Lambda)$, which is naturally a closed subscheme of $\mathbf{G}_{\Lambda}$. Recalling the zero scheme description (5) and the regularity of the section defining $\mathrm{F}$, Lemma˜2.3 implies that the morphism
@@
is flat. In particular, $\mathrm{F}(\mathfrak{X}/\Lambda)$ can be thought of as an infinitesimal deformation of $\mathrm{F}$ in $\mathbf{G}$ over $\Lambda$. For any morphism of local Artin rings $\Lambda\rightarrow\Lambda^{{}^{\prime}}$, we have
@@
$\mathcal{H}_{\mathrm{X}/\mathbf{P}}$$\mathcal{D}_{\mathrm{X}}$$\mathcal{H}_{\mathrm{F}/\mathbf{G}}$$\mathcal{D}_{\mathrm{F}}.$$\scriptstyle{\eta_{\mathcal{H}}}$$\scriptstyle{\eta}$ (14)
The proof of our main theorem requires an analogue of Lemma˜4.1 for $\mathrm{F}\subset\mathbf{G}$.
@@
(i) By Corollary˜2.1, the spectral sequence
@@
Here $\mathrm{H}^{0}(\mathrm{S}^{3}\mathcal{S}\otimes\mathrm{S}^{3}\mathcal{S}^{\vee})$ is one-dimensional (Proposition˜2.1), and it remains to combine this with the exact sequence in cohomology associated to
$0\rightarrow\mathcal{J}_{\mathrm{F}/\mathbf{G}}\otimes\mathrm{S}^{3}\mathcal{S}^{\vee}\rightarrow\mathrm{S}^{3}\mathcal{S}^{\vee}\rightarrow\mathrm{S}^{3}\mathcal{S}|_{\mathrm{F}}^{\vee}\rightarrow 0.$
(ii) Similarly, by Corollary˜2.1 the spectral sequence
@@
We now apply Lemma˜2.5 to $\mathcal{F}=\mathrm{S}^{3}\mathcal{S}^{\vee}$ on $\mathrm{S}=\mathbf{G}$. Let $\pi:\mathcal{Z}\rightarrow\mathrm{H}^{0}(\mathrm{S}^{3}\mathcal{S}^{\vee})$ be as in Lemma˜2.5, and put $\phi=\sigma^{-1}\circ\pi$, where $\sigma$ is the isomorphism of Remark˜2.1.
@@
In other words, the deformation of $\mathrm{F}$ induced by $\phi$ is complete at $f\in\mathrm{H}^{0}(\mathcal{O}_{\mathbf{P}}(3))$. For $d\geqslant 6$, this is the content of Theorem 5.3 of [5].
For a point $x_{0}$ of $\mathrm{X}$ we let $\Sigma_{x_{0}}\subset\mathbf{G}$ be the subscheme parametrising lines $\mathrm{L}$ containing $x_{0}$, and define $\mathrm{F}_{x_{0}}=\mathrm{F}\cap\Sigma_{x_{0}}$. We have $\mathrm{F}_{x_{0}}\neq\varnothing$ since $\mathrm{X}$ can be covered by lines; in fact,
$\dim\mathrm{F}_{x_{0}}\geqslant d-4$ (15)
as $\Sigma_{x_{0}}$ is a projective space of dimension $d$ and $\mathrm{F}_{x_{0}}=\mathrm{Z}(\sigma_{f}|_{\Sigma_{x_{0}}})$.
@@
For $d\geqslant 5$ there is a canonical isomorphism
@@
$\alpha:\operatorname{Aut}(\mathrm{X})\rightarrow\operatorname{Aut}(\mathrm{F}).$ (16)
If $\phi:\mathrm{X}\rightarrow\mathrm{X}$ is an automorphism, then $\alpha(\phi):\mathrm{F}\rightarrow\mathrm{F}$ satisfies
$\alpha(\phi)([\mathrm{L}])=[\phi(\mathrm{L})]$
for every line $\mathrm{L}\subset\mathrm{X}$. We now show that $\alpha$ is injective; let $\phi$ be in the kernel of $\alpha$. By (15) $\mathrm{X}$ is covered by lines, and so it suffices to show that $\phi|_{\mathrm{L}}:\mathrm{L}\rightarrow\phi(\mathrm{L})=\mathrm{L}$ is the identity for every line $\mathrm{L}\subset\mathrm{X}$. For a point $x_{0}$ of $\mathrm{L}$, $\phi(x_{0})$ lies in $\phi(\mathrm{M})=\mathrm{M}$ for every line $\mathrm{M}\subset\mathrm{X}$ containing $x_{0}$. Since the space $\mathrm{F}_{x_{0}}$ of such lines $\mathrm{M}$ has dimension $\geqslant 1$ by (15), this is possible only if $\phi(x_{0})=x_{0}$. Proposition 4 of [7] shows that the image of (16) is the subgroup $\operatorname{Aut}(\mathrm{F},\mathcal{O}_{\mathrm{F}}(1))$ of automorphisms of $\mathrm{F}$ preserving the Plücker polarization. Since $\mathrm{H}^{1}(\mathcal{O}_{\mathrm{F}})=0$, $\mathrm{H}^{0}(\Theta_{\mathrm{F}})$ is the tangent space of $\operatorname{Aut}(\mathrm{F},\mathcal{O}_{\mathrm{F}}(1))$ at the identity and so the differential of $\alpha$ at the identity gives the desired isomorphism $\mathrm{H}^{0}(\Theta_{\mathrm{X}})\xrightarrow{\sim}\mathrm{H}^{0}(\Theta_{\mathrm{F}})$. ∎
@@
$\mathrm{H}^{0}(\mathcal{N}_{\mathrm{X}/\mathbf{P}})$$\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{X}},\mathcal{O}_{\mathrm{X}})$$\mathrm{H}^{0}(\mathcal{N}_{\mathrm{F}/\mathbf{G}})$$\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$$\scriptstyle{d\eta_{\mathcal{H}}}$$\scriptstyle{d\eta}$
of differentials induced by (14). By Lemma˜2.4 (i) and Lemma˜5.1 (ii), the differential
@@
by (5). Since $\sigma$ is an isomorphism and the restriction map
@@
is surjective by Lemma˜5.1 (i), it follows that $d\eta_{\mathcal{H}}$ is surjective. It now suffices to show that
@@
associated to the conormal sequences of $\mathrm{X}\subset\mathbf{P}$ and $\mathrm{F}\subset\mathbf{G}$, respectively. By Lemma˜5.2 we have $h^{0}(\Theta_{\mathrm{X}})=h^{0}(\Theta_{\mathrm{F}})$, while
@@
result from Lemma˜4.1, Lemma˜5.1, and Example˜2.1.
If $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$, then $\mathrm{H}^{0}(\Theta_{\mathrm{F}})=0$ by Lemma˜5.2. Hence both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ are pro-representable; since $\mathcal{D}_{\mathrm{X}}$ is smooth and $d\eta$ bijective, it remains to apply Lemma˜2.2 (ii). ∎
@@
This is a consequence of the proof of Theorem˜5.1 rather than Theorem˜5.1 itself. The proof shows that $d\eta_{\mathcal{H}}$ can be identified with the isomorphism
@@
induced by $\sigma$. As $\mathcal{H}_{\mathrm{X}/\mathbf{P}}$ and $\mathcal{H}_{\mathrm{F}/\mathbf{G}}$ are pro-representable, and $\mathcal{H}_{\mathrm{X}/\mathbf{P}}$ smooth, $\eta_{\mathcal{H}}$ is an isomorphism by Lemma˜2.2 (ii). Finally, $\eta$ is surjective by Lemma˜2.2 (i), as both $\mathcal{D}_{\mathrm{X}}$ and $\mathcal{D}_{\mathrm{F}}$ have a pro-representable hull by Schlessinger’s theorem, Theorem˜2.3 (i). ∎
@@
Through the proof of Lemma˜5.2, our proof of Theorem˜5.1 depends on the results of [7]. One could get rid of this dependence by establishing a commutative diagram
@@
where the isomorphism on the left is induced by Chow’s isomorphism $\operatorname{Aut}(\mathbf{P})\rightarrow\operatorname{Aut}(\mathbf{G})$, and Lemma˜4.1 (ii), Lemma˜5.1 (ii). We expect $\eta$ to be an isomorphism without assuming the condition $\mathrm{H}^{0}(\Theta_{\mathrm{X}})=0$.
@@
There are a number of follow-up questions. If $\mathrm{X}$ is a Lefschetz cubic with a node at $x_{0}$, then the singular locus of $\mathrm{F}$ can be identified with a smooth complete intersection $\Sigma\subset\mathbf{P}_{d}$ of type $(2,3)$. The scheme $\mathrm{F}$ has rational singularities, and the blow-up
@@
of $\mathrm{F}$ along $\Sigma$ provides a resolution of singularities of $\mathrm{F}$ [8, Theorem 7.8]. In such a situation, a general construction of Wahl [27, Remark 1.4.1] yields a blow-down morphism
@@
Here $\tilde{\mathrm{F}}$ is closely related to the Hilbert scheme of points $\Sigma^{[2]}$. By [3, Theorem 36], one has a canonical isomorphism $\mathrm{H}^{1}(\Theta_{\Sigma})\xrightarrow{\sim}\mathrm{H}^{1}(\Theta_{\Sigma^{[2]}})$, which shows in particular that $\mathrm{H}^{1}(\Theta_{\Sigma^{[2]}})$ has dimension $\binom{d+2}{3}$; since this is also the dimension of $\operatorname{Ext}^{1}(\Omega^{1}_{\mathrm{F}},\mathcal{O}_{\mathrm{F}})$, the morphism $\beta$ might be an isomorphism.